# Deletion-correcting codes for an adversarial nanopore channel

### v1

@@ III Explicit construction of codes
Since $t\leq(\ell+1)/2$, by Theorem 3 two adjacent check patterns with different period $T_{1},T_{2}$ can overlap in at most $T_{1}+T_{2}-\gcd(T_{1},T_{2})-1\leq 2t-3\leq\ell-2$ positions. Moreover, if $T_{1}=T_{2}$, then by Theorem 3 they can overlap in at most $T_{1}-1\leq t-1<\ell-2$ position. Hence, any two adjacent check patterns in a string can overlap in at most $\ell-2$ positions. ∎
@@
Assume $t\leq\ell-2$. At most $\tau\leq t$ check patterns in $\tilde{\mathbf{y}}$ can have length smaller than their corresponding check patterns in $\tilde{\mathbf{x}}$.
@@
Since there are $\tau$ deletions in the output $\ell$-mers, there are $\tau$ deleted symbols in $\tilde{\mathbf{y}}$. Note that bursts are by definition disjoint. It follows that there can be at most $\tau$ deletion bursts in $\tilde{\mathbf{y}}$. So it suffices to show that a single deletion burst can reduce the length of at most one check pattern in $\tilde{\mathbf{x}}$. By Lemma 4, a deletion burst does not create any inconsistency if and only if it lies within a sufficiently long periodic pattern and the number of deletions in the burst is a multiple of the least period of that pattern. Now suppose a burst of $b\leq\tau$ deletions lies within the overlap between two sufficiently long check patterns in $\tilde{\mathbf{x}}$ with least period $T_{1}$ and $T_{2}$. Then the burst must preserve consistency with respect to both patterns, so $\operatorname{lcm}(T_{1},T_{2})$ must divide $b$. It follows that $b\geq\operatorname{lcm}(T_{1},T_{2})$. At the same time, $b$ cannot exceed the length of the overlap between the two check patterns. Therefore, by Theorem 3, we further obtain $\operatorname{lcm}(T_{1},T_{2})\leq T_{1}+T_{2}-\gcd(T_{1},T_{2})-1$. Let $g=\gcd(T_{1},T_{2})$ and write $T_{1}=a_{1}g,T_{2}=a_{2}g$. Then
$\operatorname{lcm}(T_{1},T_{2})-(T_{1}+T_{2}-\gcd(T_{1},T_{2})-1)$ $=a_{1}a_{2}g-a_{1}g-a_{2}g+g+1$
$=(a_{1}-1)(a_{2}-1)g+1$
$>0,$
leading to a contradiction. Hence, a deletion burst cannot lie within the overlap between two sufficiently long check patterns. Lastly, consider a burst of $b$ deletions that lies in one check pattern with least period $T_{1}$ but overlaps another check pattern with least period $T_{2}$ in $b^{\prime}<b$ positions. To reduce the length of both check patterns, it is necessary that the first $b_{1}\geq b-b^{\prime}$ deletions are a multiple of $T_{1}$ while the remaining $b_{2}:=b-b_{1}\leq b^{\prime}$ deletions are a multiple of $T_{2}$. But then $T_{1}\mid b$ and $T_{1}\mid b_{1}$ imply that $b_{2}$ is also a multiple of $T_{1}$. Thus, we obtain $b^{\prime}\geq b_{2}\geq\operatorname{lcm}(T_{1},T_{2})$ and the above argument still applies by replacing $b$ by $b^{\prime}$. Therefore, no single burst can reduce the lengths of two check patterns simultaneously. ∎

### v2

# arXiv:2601.21236v2[cs.IT] 02 Mar 2026

@@ III Explicit construction of codes
Since $t\leq(\ell+1)/2$, by Theorem 3 two adjacent check patterns with different period $T_{1},T_{2}$ can overlap in at most $T_{1}+T_{2}-\gcd(T_{1},T_{2})-1\leq 2t-3\leq\ell-2$ positions. Moreover, if $T_{1}=T_{2}$, then by Theorem 3 they can overlap in at most $T_{1}-1\leq t-1<\ell-2$ positions. Hence, any two adjacent check patterns in a string can overlap in at most $\ell-2$ positions. ∎
@@
Assume $t\leq\min\{\ell-2,(\ell+1)/2\}$. At most $\tau\leq t$ check patterns in $\tilde{\mathbf{y}}$ can have length smaller than their corresponding check patterns in $\tilde{\mathbf{x}}$.
@@
By Lemma 6, there is a one-to-one correspondence between check patterns in the input string $\tilde{\mathbf{x}}$ and the output string $\tilde{\mathbf{y}}$ that matches each check pattern in $\tilde{\mathbf{x}}$ with the same pattern (possibly with reduced length) in $\tilde{\mathbf{y}}$. Since there are $\tau$ deletions in the output $\ell$-mers, there are $\tau$ deleted symbols in the output string $\tilde{\mathbf{y}}$. Bursts are by definition disjoint, so there can be at most $\tau$ deletion bursts in $\tilde{\mathbf{y}}$, with equality if and only if each burst has length one. We next show that a single deletion burst can reduce the length of at most one check pattern in the input string $\tilde{\mathbf{x}}$.
Consider a deletion burst of length $b\leq\tau$ that deletes the symbols $(x_{i+1},\ldots,x_{i+b})$. By Lemma 4, the output $\ell$-mers remain consistent if and only if $(x_{i-\ell+2},\dots,x_{i+b})$ is a periodic pattern with least period $T$ satisfying $T\mid b$. Let $\mathbf{p}_{1}$ be the maximal periodic pattern in $\tilde{\mathbf{x}}$ with least period $T$ that contains $(x_{i-\ell+2},\dots,x_{i+b})$. It follows that $\mathbf{p}_{1}$ is a check pattern.
Suppose the same burst also shortens the length of another check pattern $\mathbf{p}_{2}$. Then the burst must overlap $\mathbf{p}_{2}$, and thus $\mathbf{p}_{1}$ overlaps $\mathbf{p}_{2}$. By Lemma 5, the size of the overlap between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ is at most $\ell-2$. Since check patterns have length at least $\ell-1$, it follows that neither $\mathbf{p}_{1}$ nor $\mathbf{p}_{2}$ is contained in the other, and thus the overlap between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ must be a suffix of one and a prefix of the other. Since the length of $\mathbf{p}_{1}$ is at least $b+\ell-1$, it follows that $\mathbf{p}_{1}$ contains at least $b+\ell-1-(\ell-2)>b$ consecutive symbols that are disjoint from $\mathbf{p}_{2}$.
Since $\mathbf{p}_{1}$ has period $T$ and $T\mid b$, deleting any $b$ consecutive symbols in $\mathbf{p}_{1}$ produces the same resulting string as deleting $(x_{i+1},\ldots,x_{i+b})$. Therefore, one may assume without loss of generality that the burst deletes any $b$ consecutive symbols of $\mathbf{p}_{1}$. In particular, one may replace the original burst by an equivalent burst of length $b$ that is disjoint from $\mathbf{p}_{2}$, without changing the output string. Under this equivalent burst, $\mathbf{p}_{2}$ is unaffected and cannot be shortened by this burst, which is a contradiction.
Therefore, a single deletion burst can shorten at most one check pattern. Since there are at most $\tau$ bursts, at most $\tau$ check patterns in $\tilde{\mathbf{y}}$ can have length smaller than their corresponding check patterns in $\tilde{\mathbf{x}}$. ∎